What is the mass of a sample of Aluminum that absorbs 596 Joules of heat when the temperature is raised from 30.0 degrees C to 100.1 degrees C?
1
Expert's answer
2019-05-14T03:40:51-0400

We will use this formula, to count out the mass of Aluminium:

Q = cmΔT = 596 J

The specific heat capacity (c) for Aluminium is:

c = 0.91 kJ/ kg * K, so we have to do a convertation.

Q = 596 J = 0.596 kJ;

ΔT = 100.1 C - 30.0 = 70.1 C + 273.15 = 343.25 K;

Now we can calculate the mass of Aluminium:

m = Q/cΔT = 0.596 / (0.91*343.25) = 0.596 / 312.3575 = 0.0019 kg = 1.9 g.